OpenStudy (anonymous):

you hav already walked 5 miles of an 18 mile trail. if you walk the rest of the trail at a pace of 1 mile in 17 minutes, how many hours will it take you to finish trail? round the answer to the nearest tenth

OpenStudy (anonymous):

t= 17(18-5) t being time t=17(13) t=221 minutes h=221/60 h=3.683333333 repeating now round to 3.7 hours